OpenStudy (jiteshmeghwal9):

If \(\Large{a+b+c=0}\) the value of \(\Large{\frac{a^2}{bc}+\frac{b^2}{ca}+\frac{c^2}{ab}}\) is :-

hartnn (hartnn):

know the formula for a^3+b^3+c^3 =... ?

OpenStudy (jiteshmeghwal9):

3abc

hartnn (hartnn):

hint : first term = a^2/bc = a^3/ abc

OpenStudy (jiteshmeghwal9):

what's the use of this hint @hartnn ??

OpenStudy (jiteshmeghwal9):

ohhh ! gt it

hartnn (hartnn):

your expression = (a^3+b^3+c^3/abc ) = (a++b+c)(....)/abc good!

OpenStudy (jiteshmeghwal9):

we should use this hint in every frac & make the denominators same & solve

hartnn (hartnn):

mostly this works....

OpenStudy (jiteshmeghwal9):

hmm... thanx a lot @hartnn:)